OpenStudy (maheshmeghwal9):

The initial velocity of a body thrown at an angle to the horizon is \[V_0\]The maximum range is S. At what angle \[\alpha \]to the horizon should the body be thrown to have a range equal to \[l \space ? \space (l

OpenStudy (anonymous):

So you just want an angle that is less than S? To know this, you should know that 45degrees is the maximum angle for any launch.

OpenStudy (maheshmeghwal9):

This is nt the answer & i think here this approach isn't workable.

OpenStudy (maheshmeghwal9):

& one more thing @Coda 45 degrees is nt the maximum angle It is the angle at which the range is maximum.

OpenStudy (anonymous):

You want a range that is less than the maximum range. (l<s) Knowing the maximum range is achieved by 45degrees. What does that tell you?

OpenStudy (maheshmeghwal9):

I don't know but the answer of this question in my book is given as: -\[\LARGE{\color{red}{\alpha = \frac{1}{2} \sin^{-1} \frac{gl}{V_0^2}.}}\]

OpenStudy (anonymous):

Alright, where do you think you start using the kinematics / mechanics?

OpenStudy (maheshmeghwal9):

wt do u mean?

OpenStudy (anonymous):

Sorry, I meant to say what is it exactly you need help with? I can draw a diagram and we can go from there if you would like.

OpenStudy (maheshmeghwal9):

ya i would like but how to get that answer which i gave from my book!

OpenStudy (vincent-lyon.fr):

Work out range for fixed initial speed, then solve for angle. There should be 2 solutions for each value of range less than maximum range.

OpenStudy (maheshmeghwal9):

u mean like this. \[S=\frac{V_0^2 \sin2( 45^o) }{g}=\frac{V_0^2}{g}\] \[l<S.\]Therefore \[l=\frac{V_0^2 \sin2 \alpha}{g}<\frac{V_0^2}{g}\] But now wt to do @Vincent-Lyon.Fr sir ?

OpenStudy (maheshmeghwal9):

Actually I am nt getting the answer:(

OpenStudy (vincent-lyon.fr):

You have worked out: \(l=\Large \frac{V_0^2 \sin2\alpha }{g}\), which means \(\sin2\alpha=\Large \frac {gl}{V_0^2}\) for \(l<\Large \frac {V_0^2}{g}\) the sine has two solutions: \(2\alpha=\sin^{-1}\Large \frac {gl}{V_0^2}\) or \(2\alpha=\pi-\sin^{-1}\Large \frac {gl}{V_0^2}\) which leads to: \(\alpha_1=\Large \frac12\sin^{-1}\Large \frac {gl}{V_0^2}\) or \(\alpha_2=\Large \frac\pi2- \frac12\sin^{-1}\Large \frac {gl}{V_0^2}\) Shot for \(\alpha_1\) is called 'straight shot', as opposed to \(\alpha_2\). I do not know the English term for it, but the French call it 'bell shot'. It might be the same in English.

OpenStudy (maheshmeghwal9):

oh thanx a lot sir:)

OpenStudy (maheshmeghwal9):

gt it ^_^